OpenStudy (anonymous):

A line makes angles A, B,C,D with the four diagonals of a cube show that Cos^2A+cos^2B+cos^2C+cos^2c+cos^2D=4/3

OpenStudy (paxpolaris):



OpenStudy (anonymous):

What is this I can't understand

OpenStudy (paxpolaris):

do you have a way of doing this... and are just stuck at some point? i think i could prove this using vectors ... but would like to know what method you might be trying?

OpenStudy (paxpolaris):

let 2 of the diagonals be in the XY plane an the other two in the YZ planes. These two planes intersect at the Y-axis. and all 4 diagonals meet at the origin. in the XY-plane the diagonals are: \[y=\frac1{\sqrt2}x\]\[y=-\frac1{\sqrt2}x\] and similarly in the YZ plane, the diagonals are \[y=\frac1{\sqrt2}z\]\[y=-\frac1{\sqrt2}z\]

OpenStudy (paxpolaris):

Since all lines pass through the origin, we can represent them using Position Vectors of points on those lines. in the XY plane, we can replace the line \(y={1\over \sqrt2}x\) with the vector \(\left< \sqrt2,1,0\right>\). Let the unknown line be represented by the UNIT-vector \(\left< x,y,z\right>\) ... {it's magnitude is 1} \(A\) is the angle between \(\left< x,y,z\right>\) and \(\left< \sqrt2,1,0\right>\). \[\large \cos A={\sqrt2\cdot x +y \over 1\cdot \sqrt3}\]\[\large \cos^2 A={2x^2 +y^2+2\sqrt2y \over 3}\]

OpenStudy (paxpolaris):

\(B\) is the angle between \(\left< x,y,z\right>\) and \(\left< \sqrt2,-1,0\right>\) .... { the diagonal \(y=-x/\sqrt2 \)} \[\large \cos^2 B={2x^2 +y^2-2\sqrt2y \over 3}\] \[\therefore\cos^2A+\cos^2B= {4x^2+2y^2 \over 3}\]

OpenStudy (paxpolaris):

\(C\) is the angle between \(\left< x,y,z\right>\) and \(\left< 0,1,\sqrt2\right>\) .... { the diagonal \(y=z/\sqrt2\) } \[\large \cos^2 C={2z^2 +y^2+2\sqrt2yz \over 3}\] \(D\) is the angle between \(\left< x,y,z\right>\) and \(\left< 0,-1,\sqrt2\right>\) .... { the diagonal \(y=-z/\sqrt2\) } \[\large \cos^2 D={2z^2 +y^2-2\sqrt2yz \over 3}\] \[\therefore\cos^2C+\cos^2D= {4z^2+2y^2 \over 3}\]

OpenStudy (paxpolaris):

\[\cos^2A+\cos^2B+\cos^2C+\cos^2D\\= {4x^2+4y^2+4z^2\over 3}\\\\= {4\left( x^2+y^2+z^2 \right) \over 3}\\=\frac43\] because by definition of unit vector \(x^2+y^2+z^2=1\)